# TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval

### v2

@@ 3.2. Opportunities of GPU-accelerated Retrieval with Limited Memory
(Gao et al., 2023a) (Liu, 2022) (Liu, 2022) (Shao et al., 2023) (Jiang et al., 2023c) (Asai et al., 2023)
NQ (Kwiatkowski et al., 2019) 73.1% 63.2% 91.5% 83.8% 79.1% 100.0%
HotpotQA (Yang et al., 2018) 75.3% 62.5% 89.6% 89.4% 80.2% 100.0%
TriviaQA (Joshi et al., 2017) 73.1% 61.6% 86.2% 86.1% 81.7% 100.0%
Table 1. IVF cluster overlapping rate between the input and output of the pre-retrieval generation. The nprobe is set to 512. Since Self-RAG does not incorporate query transform, its coverage is always $100\%$.
3.3. Overlapping of IVF Clusters
@@
Similarity of queries at different stages. During the pre-retrieval process of RAG pipelines (e.g., query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022)), an LLM call is issued to refine an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$, which is then used for retrieval. This process often rewrites the query into a different format (Gao et al., 2023a) or simplifies its complexities (Liu, 2022), which intuitively preserves the query’s core semantic content. Hence, the embedding vectors of $q_{\text{in}}$ and $q_{\text{out}}$ are likely to be similar. This similarity, in turn, suggests that the IVF clusters to which these queries would be assigned will overlap. Therefore, $q_{\text{in}}$ can serve as a valuable hint for predicting subsequent memory accesses.
@@
Prediction coverage. To verify this hypothesis, we evaluate the average cluster coverage rate between prefetched clusters and clusters used for retrieval in three popular question-answer datasets (NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017)) and six RAG pipelines. Table 1 shows the coverage when we prefetch $256$ clusters. From the table, we observe that IVF cluster overlap rates are consistently high across a range of datasets and pipelines. For instance, even the lowest reported values remain above 61.6% (for SubQ).
@@
Opportunity. This data shows an opportunity for predicting required clusters, which can make it possible to hide data transfer overhead during LLM generation. In this paper, we aim to leverage this observation to accelerate the inference latency for RAG.
Figure 6. (a) The overview of lookahead retrieval compared against CPU-offloaded retrieval, and (b) the system design of TeleRAG. After identifying clusters ($C_{\text{in}}$) for the initial query $q_{\text{in}}$, $C_{\text{in}}$ is transferred to the GPU while concurrently generating $q_{\text{out}}$. At retrieval time, the GPU searches prefetched clusters while the CPU handles remaining ones, before results are merged.
4. Design of TeleRAG
Based on the high IVF cluster overlapping observation presented in §3.3, we propose TeleRAG, an efficient RAG inference system that utilizes lookahead retrieval to prefetch a minimal set of likely IVF clusters to GPU, and accelerate the retrieval process. In this section, we detail the design of lookahead retrieval and implementation of TeleRAG.
4.1. Overview
@@
Figure 6 presents the overview of TeleRAG with lookahead retrieval. In Figure 6, $q_{\text{in}}$ denotes the initial input to the pre-retrieval stage and $q_{\text{out}}$ denotes its output, i.e., the query given to the subsequent retrieval stage. We highlight the IVF clusters corresponding to $q_{\text{in}}$ and $q_{\text{out}}$ with a red background ($C_{\text{in}}$), and a blue background ($C_{\text{out}}$), respectively. Due to the semantic similarity of $q_{\text{in}}$ and $q_{\text{out}}$, there is significant overlap between $C_{\text{in}}$ and $C_{\text{out}}$, marked with a purple background ($C_{\text{overlap}}$).
Given this, the proposed lookahead retrieval technique in TeleRAG operates in three key steps:
  1. (1)
    Pre-retrieval & data transfer: During LLM generation, identify $C_{\text{in}}$ and transfer corresponding data to GPU memory, prioritizing clusters closest to $q_{\text{in}}$. This leverages GPU DMA for concurrent transfer of clusters data and compute of LLM generation.
  2. (2)
@@
    GPU similarity search: After identifying $C_{\text{out}}$, the GPU efficiently performs a similarity search on the predicted clusters ($C_{\text{overlap}}$) already in its memory.
  3. (3)
@@
    CPU similarity search: Concurrently with the GPU search, the CPU processes the similarity search for the missed clusters ($C_{\text{miss}}$) that were not prefetched to the GPU, reducing the CPU’s overall workload.
In summary, our lookahead retrieval speeds up retrieval by overlapping data prefetching to the GPU with LLM generation and performing concurrent similarity searches on both the GPU (for predicted clusters) and CPU (for the remainder), significantly reducing the CPU’s computational workload.
4.2. Finding the Optimal Prefetch Amount
(1) $t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$
As shown in §3.1, GPU retrieval ($t_{g}$) is generally much faster than CPU retrieval ($t_{c}$). Thus, we assume $t_{c}\gg t_{g}$. As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed (Jiang et al., 2023a):
(2) $t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$
We now examine two cases:
    (3) $b_{p}^{*}=B\cdot t_{\text{LLM}}.$
    (4) $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$
    From this, we obtain:
    (5) $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$
In summary, our analysis shows that the optimal prefetch amount $b_{p}^{*}$ can only lie at one of two points: (1) Prefetch until LLM generation completes (i.e. $b_{p}=B\cdot t_{\text{LLM}}$). (2) A point determined by Eq. 5. However, under typical CPU-GPU bandwidth (e.g., 55 GB/s on PCIe 5), the time spent loading additional clusters often outweighs any retrieval latency reduction from lowering $r_{\mathrm{miss}}$. Consequently, the second scenario in Eq. 5 becomes nearly infeasible in practice. Therefore, on current hardware, prefetching exactly until LLM execution ends is generally the most effective choice.
Profiling-guided approach. Although $b_{p}^{*}$ can be derived from the analysis above, it depends on knowing $t_{\mathrm{LLM}}$ for each query, which cannot be obtained ahead of time. To address this, we use a profiling-guided approach, leveraging the observation that, despite differences in query content, the output length (and thus generation time) for a RAG pipeline often remains similar across most queries. Accordingly, for each RAG pipeline, we can measure $t_{\mathrm{LLM}}$ on a calibration set containing n queries, and get estimated $\hat{b}_{p}^{*}=B\cdot\overline{t}_{\mathrm{LLM}}$, where $\overline{t}_{\mathrm{LLM}}=\mathrm{mean}\{t_{\mathrm{LLM},1},t_{\mathrm{LLM},2},\dots,t_{\mathrm{LLM},n}\}.$ This estimated $\hat{b}_{p}^{*}$ can then be used for incoming queries, ensuring a near-optimal prefetch amount.
4.3. Design Details
@@
Sorting Optimization. TeleRAG accelerates the final $k$-argmin sorting step of IVF search by leveraging GPU. Because each cluster may contain thousands of points, sorting is particularly well-suited to GPU parallelism (Arkhipov et al., 2017). To enable this, TeleRAG transfers the scalar distance values computed on the CPU for $C_{\text{miss}}$ to the GPU. Unlike full vector data, this transfer incurs negligible overhead. The GPU then performs a global sort over the combined distances from $C_{\text{overlap}}$ and $C_{\text{miss}}$, delivering end-to-end speedups than the CPU-only retrieval systems.
Prefetch Target. TeleRAG’s prefetching mechanism targets a specific number of bytes ($b_{p}$) rather than the number of clusters to ensure more predictable performance. Since IVF cluster sizes are often highly uneven, targeting a fixed cluster count can lead to unstable data loading times, whereas a byte limit provides a clearer upper bound on the transfer duration based on bandwidth.
When selecting clusters to meet the prefetch byte budget ($b_{p}$), the system adds whole clusters sequentially based on proximity to the query. If adding the next closest cluster would exceed the budget, it is omitted entirely rather than being partially loaded, ensuring a clean division where each cluster is processed exclusively by either GPU or CPU.
@@
Multi-Round Support. For multi-round RAG involving the same input query, the system performs a full prefetch (up to the budget) only in the first round, leveraging high cluster similarity across rounds. In subsequent rounds, it incrementally fetches only the additional required clusters that were not loaded previously, optimizing data transfer while respecting the memory budget.
Batch Support. For batch inference, a challenge for lookahead retrieval is that each query requires a different set of IVF clusters. In our design, we fix the prefetching budget to the amount we got from §4.2 irrespective of batch size, and equally distribute the budget to each query in a batch. As a result, although the average cluster hit rate per request will decrease for larger batch sizes, we can ensure each query’s retrieval search time is equally accelerated, forming a balanced batched retrieval.
Multi-GPU support. For multi-GPU systems, we perform a greedy search based on embedding similarities of queries to group similar queries into mini-batches, and distribute the mini-batches among GPUs. The greedy search is efficient and incurs negligible overhead: In our profiling for up to the batch size of 256, the latency to perform greedy search is less than $0.1$ s, which is minimal compared to the end-to-end latency of each batch.
4.4. Implementation Details
TeleRAG is implemented in Python and also leverages PyTorch’s (Ansel et al., 2024) operator ecosystem for efficient computation. The datastore index is initially constructed using FAISS (Douze et al., 2024), and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors.
To enable concurrent CPU-GPU data transferring and LLM generation, we utilize the _copy(non_blocking=True) API in PyTorch and use separate CUDA streams for prefetching and LLM. In this way, the data copy operations will not block the GPU to perform computation, and thus the TeleRAG can do prefetching in parallel to the pre-retrieval LLM generation.
To implement the index search with GPU-CPU cooperation, for the GPU part, we use a single matrix-vector multiplication that computes distances for all prefetched vectors; for the CPU part, we utilize multithreading in Python to parallelize similarity searches across clusters. We then move the distances computed from CPU to GPU, merge with distances on GPU and perform global sorting on GPU.
Specification Value
Dataset wiki_dpr (Karpukhin et al., 2020)
Dataset size 2.1 billion tokens
#\# of chunks 21 million
#\# of IVF cluster 4096
Embed model Contriever (Izacard et al., 2021)
Embed dimension 768
Distance metric Inner Product
Index size 61 GB
Table 2. Detailed configurations of our retrieval index.
5. Evaluation
@@
We conduct extensive experiments to evaluate the effectiveness of TeleRAG. In this section, we describe the necessary details on how we set up the evaluations, present experimental results, and provide in-depth analysis and discussions.
5.1. Evaluation Datasets and RAG Models
@@
Datastore. We build a datastore based on the wiki_dpr dataset (Karpukhin et al., 2020), a popular dataset that contains 2.1 billion tokens from Wikipedia. Following previous works (Asai et al., 2023; Karpukhin et al., 2020; Min et al., 2023), we chunk the passages by every 100 tokens, and use Contriever (Izacard et al., 2021) to generate an embedding for each chunk. The embeddings have a hidden dimension of 768.
Vector index. For the baseline retrieval, we build an IVF vector index using Faiss (Douze et al., 2024) on the datastore. As described in §4.4, we convert the Faiss index to a customized index in PyTorch for TeleRAG. See Table 2 for the detailed configurations of our vector index and datastore.
@@
LLMs. We evaluate TeleRAG on the Llama model family (Touvron et al., 2023) in three different sizes (Llama-3.2-3B, Llama-3-8B, Llama-2-13B) to represent different use cases.
Setup Desktop Server
CPU Threadripper 5975 EPYC 9554
CPU memory size 512 GB 1.5 TB
GPU RTX4090 (NVIDIA, 2024a) H100 (NVIDIA, 2024b)
GPU memory size 24 GB 80 GB
CPU-GPU Bus PCIe 4 PCIe 5
Bandwidth n 32 (24) GB/s n n 64 (51) GB/s n
Table 3. Hardware specifications for our setups. In bandwidth, the number in the parentheses is the actual bandwidth we measured from our system.
@@
RAG pipelines. We evaluate TeleRAG with six popular RAG pipelines. Figure 7 shows the overview of each pipeline. Note that although some pipelines lack pre-retrieval generation, the post-retrieval generation serves similar functionality for the retrieval of the next iteration. Below are brief descriptions of the evaluated RAG pipelines.
  1. (1)
@@
    HyDE (Gao et al., 2023a) prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
  2. (2)
@@
    SubQestion (SubQ) [^3] Implemented in LlamaIndex as SubQuestionQueryEngine. (Liu, 2022) prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
  3. (3)
@@
    Iterative (Iter) [^4] Implemented in LlamaIndex as MultiStepQueryEngine. (Liu, 2022) prompts LLM to generate narrower questions first and iteratively refine them based on previous answers. At the end of each iteration, it prompts LLM to judge if the answer is good enough.
  4. (4)
@@
    Iter-RetGen (IRG) (Shao et al., 2023) iteratively do retrieval and LLM generation for 3 iterations.
  5. (5)
@@
    FLARE (Jiang et al., 2023c) iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
  6. (6)
@@
    Self-RAG (S-RAG) [^5] We evaluate on the short-form version which only has one iteration. (Asai et al., 2023) uses the LLM to judge for retrieval, generate responses, and self-critique on the responses. We use fine-tuned models based on Llama2-7B and Llama2-13B from their official repository (Asai et al., 2024b).
Figure 7. Overview of six RAG pipelines that we evaluate.
@@
Evaluation Datasets. We use three commonly used question-answering datasets, NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017). For each dataset, we randomly sample 1024 queries and report the average unless otherwise specified.
5.2. Experiment Setups

### v3

@@ 3.2 GPU-accelerated Retrieval with Runtime Transfer
NQ 73.1% 63.2% 91.5% 83.8% 79.1% 100.0%
HotpotQA 75.3% 62.5% 89.6% 89.4% 80.2% 100.0%
TriviaQA 73.1% 61.6% 86.2% 86.1% 81.7% 100.0%
Table 1: IVF cluster overlapping rate between the input and output of the pre-retrieval generation. Since Self-RAG does not incorporate query transform, its coverage is always $100\%$.
Figure 6: The overview of TeleRAG system. When queries arrive, some clusters are already cached on the GPU from previous requests. These queries are grouped into micro-batches based on semantic similarity. For each micro-batch, TeleRAG prefetches the required clusters using lookahead retrieval in parallel with the pre-retrieval generation. During retrieval, TeleRAG performs a hybrid search: the GPU processes cache hits directly from GPU memory, while the CPU handles cache misses. The retrieved documents are then passed to the post-retrieval generation stage.
3.3 Overlapping of IVF Clusters
@@
Similarity of queries at different stages. During RAG pre-retrieval stages (e.g., query transformation gao2022precise_hyde; zheng2023take; peng2023_beque; openscholar), an LLM refines an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$ for the actual retrieval. While this process often rewrites or simplifies the query, it intuitively preserves its core semantic content. Consequently, the embedding vectors of $q_{\text{in}}$ and $q_{\text{out}}$ are likely to be similar, suggesting their target IVF clusters will significantly overlap. Therefore, $q_{\text{in}}$ can serve as a valuable hint for predicting $q_{\text{out}}$.
@@
Prediction coverage. To verify this hypothesis, we evaluated the average cluster coverage (predicted vs. actual) in three popular QA datasets (NQ kwiatkowski2019natural, HotpotQA yang2018hotpotqa, and TriviaQA joshi2017triviaqa) and six RAG pipelines. As Table 1 shows, when prefetching 256 clusters, the overlap is consistently high. For instance, even the lowest reported coverage (for SubQ) remains above 61.6%.
@@
Opportunity. This data shows an opportunity to predict required clusters, hiding data transfer overhead during LLM generation. In this paper, we aim to leverage this observation to accelerate the inference latency for RAG.
4 Design of TeleRAG
Driven by the need for high-performance but memory-efficient RAG inference, we propose TeleRAG, an end-to-end system that accelerates retrieval through its core lookahead retrieval mechanism, which accelerates the retrieval process with a GPU but only requires a small fraction of the datastore on the GPU memory. Alongside this technique, TeleRAG further designs optimizations on caching, prefetching amount, and scheduling algorithms for batching and multi-GPU scenarios, forming a complete system for both local and cloud usages.
This section details the lookahead retrieval mechanism (§4.1), and then describes how TeleRAG extends it with batching and multi-GPU support (§4.2), followed by other system-level optimizations (§4.3). Appendix C provides implementation details.
4.1 Lookahead Retrieval Mechanism
The core mechanism of TeleRAG, lookahead retrieval, is inspired by the observation in §3.3 that queries across different RAG stages are highly correlated and therefore select overlapping IVF clusters. Building on this insight, lookahead retrieval predicts and prefetches likely-needed clusters to the GPU during LLM generation, overlapping data transfer with computation. During retrieval, TeleRAG leverages lookahead retrieval to coordinate the CPU and GPU so that prefetched data are processed on the GPU, while the CPU concurrently handles any missed clusters. This cooperative execution forms the foundation of TeleRAG’s low-latency, memory-efficient design.
@@ 4.1 Lookahead Retrieval Mechanism
Figure 6 shows the overview of TeleRAG with lookahead retrieval. Let $q_{\text{in}}$ denote the query input to the pre-retrieval stage and $q_{\text{out}}$ denote the output passed to the retrieval stage. The IVF clusters selected by $q_{\text{in}}$ and $q_{\text{out}}$ are highlighted with a green background ($C_{\text{in}}$) and a purple background ($C_{\text{out}}$), respectively. Due to the semantic similarity between $q_{\text{in}}$ and $q_{\text{out}}$, there is substantial overlap between $C_{\text{in}}$ and $C_{\text{out}}$.
Guided by this correlation, lookahead retrieval operates in the following steps:
  1. 1.
    Predict & prefetch: During LLM generation, prefetch IVF clusters likely to be used in retrieval to GPU memory, identified by their distance to $q_{\text{in}}$. This transfer is performed asynchronously via GPU DMA, overlapping with ongoing LLM computation.
  2. 2.
@@
    GPU similarity search: Once $q_{\text{out}}$ is available, the GPU efficiently searches the predicted clusters ($C_{\text{overlap}}$) already resident in GPU memory.
  3. 3.
@@
    CPU similarity search: Concurrently, the CPU performs similarity search over the remaining clusters ($C_{\text{miss}}$) that were not prefetched.
  4. 4.
    Merge: The search results from GPU and CPU are merged on GPU. The retrieval documents are fed to LLM for post-retrieval generation on GPU.
In summary, lookahead retrieval enables TeleRAG to accelerate retrieval by overlapping data prefetching with LLM generation and distributing the similarity search between GPU and CPU. This design significantly reduces CPU computation and data transfer latency, forming the backbone of TeleRAG’s efficiency.
Prefetching amount. Prefetching more clusters improves hit rate and retrieval speed, but also increases CPU–GPU transfer time. If this transfer exceeds the pre-retrieval generation window, the overlap advantage is lost and may add latency, though slight overruns are acceptable if retrieval savings are substantial. Through analytical modeling and empirical profiling on modern hardware, we found prefetching up to the pre-retrieval LLM generation time ($t_{\mathrm{LLM}}$) achieves the optimal balance between latency reduction and transfer overhead. A detailed derivation of this result is provided in Appendix B. Since $t_{\mathrm{LLM}}$ varies per query, for a given pipeline and hardware bandwidth $B_{\mathrm{link}}$, we estimate its average ($\overline{t}_{\mathrm{LLM}}$) on a calibration set and set the prefetching amount to $B_{\mathrm{link}}\times\overline{t}_{\mathrm{LLM}}$.
Figure 7: Overview of TeleRAG’s multi-GPU system design. The prefetching scheduler clusters semantically similar queries into micro-batches, while the cache-aware scheduler assigns these batches to GPUs based on cache locality.
4.2 Batching and Multi-GPU System
While §4.1 describes lookahead retrieval on a single query and single GPU basis, we now demonstrate how TeleRAG extends this mechanism to support batches and multi-GPU inference. The system’s multi-GPU architecture is illustrated in Figure 7.
Batching design. For batched scenarios, a challenge for lookahead retrieval is that each query requires a different set of IVF clusters. In our design, we apply the fixed total prefetching budget as described in §4.1, and distribute it equally among all queries in the batch. Although the per-query hit rate drops as the batch size grows, this strategy provides balanced retrieval acceleration for each query.
Multi-GPU support. In multi-GPU scenarios that serve a large number of RAG queries, TeleRAG divides a global batch of queries into micro-batches and assigns them to each available GPU. Each GPU then independently processes the full RAG pipeline for its assigned micro-batches.
Scheduling optimization for prefetching. To compensate for the reduced hit rate in larger batches, we introduce a prefetching scheduler (see Figure 7). It operates by performing a greedy search on the global batch, grouping queries by the lowest L2 distance. While this scheduling adds latency, the greedy search is highly efficient on GPUs, incurring negligible overhead. Our profiling shows that for batch sizes up to 256, the search latency is less than $0.1$ s, which is minimal compared to the end-to-end batch latency.
Cache design. To improve performance, TeleRAG implements a dynamic caching strategy for data clusters. Instead of discarding all clusters after a request is served, it retains the most frequently accessed ones in GPU memory. This allows subsequent requests to reuse these cached clusters, boosting the cache hit rate and prefetching efficiency.
Cache-aware scheduling. To further exploit the benefits of caching, we design a cache-aware scheduler (Figure 7) that assigns micro-batches to GPUs based on cache locality. The scheduler employs a greedy strategy: it first selects the micro-batch with the greatest overlap across all GPU caches and assigns it to the best-matching GPU. It then iteratively schedules the remaining micro-batches in descending order of overlap with each GPU’s cached clusters. Although this scheduling introduces minor overhead, the performance gains from improved cache locality outweigh the cost (§5.4).
4.3 System Optimizations
@@
GPU sorting. TeleRAG accelerates the final sorting stage of IVF search by leveraging the GPU. To achieve this, TeleRAG transfers the scalar distance values of $C_{\text{miss}}$ from the CPU to the GPU. Unlike transferring full vector data, this operation is lightweight and incurs negligible overhead. The GPU then performs a global sort over the combined distances of $C_{\text{miss}}$ and $C_{\text{overlap}}$.
Prefetching target. TeleRAG prefetches a fixed byte budget ($b_{p}$), rather than a cluster count, to ensure predictable transfer times given highly uneven cluster sizes. The system fills this budget by adding whole clusters sequentially based on query proximity. If the next closest cluster exceeds the remaining budget, it is skipped entirely, ensuring a clean GPU/CPU processing split.
@@
Prefetching for multi-round. For multi-round RAG involving the same input query, the system performs a full prefetch (up to the budget) only in the first round, leveraging high cluster similarity across rounds. In subsequent rounds, it incrementally fetches only the additional required clusters that were not loaded previously, optimizing data transfer while respecting the memory budget.
5 Evaluation
@@
We conducted extensive experiments to evaluate the effectiveness of TeleRAG. In this section, we describe the necessary details on evaluation setups, present experimental results, and provide in-depth analysis and discussions.
5.1 Evaluation Datasets and RAG Models
@@
Datastore. We built a datastore based on the wiki_dpr dataset karpukhin2020dense, a popular dataset that contains 2.1 billion tokens from Wikipedia. Following previous works asai2023self; karpukhin2020dense; min2023silo, we chunked the passages by every 100 tokens, and used Contriever izacard2021unsupervised to generate an embedding for each chunk. The embeddings have a hidden dimension of 768. The index size is 61 GB and we clustered the embeddings into 4096 IVF clusters. See Appendix D for detailed configurations of the index.
@@
LLMs. We evaluated TeleRAG on the Llama model family touvron2023llama in two different sizes (Llama-3.2-3B, Llama-3-8B) to represent different use cases.
Setup Desktop Server1 Server2
CPU Threadripper 5975 EPYC 9554 EPYC 9534
CPU mem. size 512 GB 1.5 TB 1.5 TB
GPU RTX4090 H100 $8\timesH$200
GPU mem. size 24 GB 80 GB 140 GB
CPU–GPU Bus PCIe 4 PCIe 5 PCIe 5
Bandwidth 32 GB/s 64 GB/s 64 GB/s
Table 2: Hardware specifications for our setups.
Figure 8: Overview of six RAG pipelines that we evaluate.
@@
RAG pipelines. We evaluated TeleRAG with six popular RAG pipelines, as depicted in Figure 8. Note that even in pipelines lacking a pre-retrieval stage, the post-retrieval generation serves a similar function for the next retrieval iteration. Below are brief descriptions of these pipelines.
  1. 1.
@@
    HyDE gao2022precise_hyde prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
  2. 2.
@@
    SubQuestion (SubQ)  sub_question_llamaindex prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
  3. 3.
@@
    Iterative (Iter)  multistep_llamaindex prompts LLM to generate narrower questions first and iteratively refine them based on previous answers. At the end of each iteration, it prompts LLM to judge if the answer is good enough.
  4. 4.
@@
    Iter-RetGen (IRG) shao2023enhancing iteratively does retrieval and LLM generation for 3 iterations.
  5. 5.
@@
    FLARE jiang2023flare iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
  6. 6.
@@
    Self-RAG (S-RAG) asai2023self uses the LLM to judge for retrieval, generate responses, and self-critique on the responses. We use the fine-tuned model based on Llama2-7B from their official repository selfrag_code for trace generation.
@@
Evaluation datasets. We use three commonly used question-answering datasets, NQ kwiatkowski2019natural, HotpotQA yang2018hotpotqa, and TriviaQA joshi2017triviaqa. For each dataset, we randomly sampled 1024 queries and reported the average. When cache is enabled, to ensure stable results and avoid an initial low hit rate from a cold start, we use 512 queries for warming up the cache and another different 512 queries for evaluation.
5.2 Experiment Setups